# Beam-splitter-free, high-rate quantum key distribution inspired by intrinsic quantum mechanical spatial randomness of entangled photons

### v1

@@ 3 Results and discussions
We first characterized EPS 1 and EPS 2 in terms of their quantum performance parameters, including entanglement visibility, Bell parameter, and state fidelity. As both sources are derived from a single entangled photon source, they exhibit nearly identical characteristics. For clarity, we present the representative results of EPS 2 (see green paths in Fig. 1(b)). The measurements, summarized in Fig. 2, were performed at a fixed pump power of 3 mW and crystal temperature of T = 41.5∘C. We recorded the coincidence counts between photons transmitted through the PBS of both Alice and Bob, with Alice having fast axis of the HWP fixed at , , , and , corresponding to the (black), (blue), (red), and (green) bases, respectively, while Bob continuously rotated the HWP from to in steps. The resulting interference fringes are shown in Fig. 2(a).
@@
As shown in Fig. 2(a), the coincidence counts exhibit the expected sinusoidal modulation, with measured fringe visibilities of 97.8, 97.8, 97.7, and 97.7 in the , , , and bases, respectively—well above the quantum threshold of 71 [31], thereby confirming high-quality entanglement. Furthermore, by analyzing coincidence counts across 16 different combinations of HWP settings of Alice and Bob and applying the standard CHSH inequality formalism [36], we obtained a Bell parameter of , clearly violating the classical bound by nearly 109 standard deviations. Further, we carried out quantum state tomography based on coincidence measurements across different polarization projections. The reconstructed real and imaginary parts of the two-qubit density matrix, shown in Fig. 2(b) and 2(c), respectively, confirm the generation of the maximally entangled Bell state with a fidelity of 97. The source brightness of EPS 2 is estimated to be 0.04 MHz/mW/nm. We observed comparable performance parameters for EPS 1. These results establish the suitability of both EPS 1 and EPS 2 for the implementation of the QKD protocol.
@@
With the quantum parameters of the entangled photon sources established, we implemented the BBM92 protocol and benchmarked its performance against the conventional beam-splitter–based approach. Keeping the crystal temperature fixed at C, we measured the sifted key rate and QBER as functions of pump power. The results are shown in Fig. 3. As expected, the sifted key rate calculated by summing all the coincidence counts of same basis detection of two users (here, Alice and Bob), increases with pump power in both cases owing to the higher pair-generation rate. In our scheme (black dots), the sifted key rate rises from 0.29 Mbps to 2.9 Mbps as the pump power is increased from 1 to 13 mW, corresponding to a slope of 0.22 Mbps/mW for a one-meter free-space link. By contrast, the conventional scheme (black squares) exhibits only a modest increase from 0.045 Mbps to 0.47 Mbps, with a slope of 0.035 Mbps/mW. Such a superior performance of our current BBM92 scheme in terms of higher sifted key rate with respect to the conventional BBM92 scheme can be understood as follows. Let us consider one of the derived sources, EPS 1. In the absence of a beam splitter, the coincidence rate between Alice and Bob in the same basis (e.g., Z) is . Introducing an ideal 50:50 beam splitter at each photon path of Alice and Bob reduces the useful coincidences to the transmitted–transmitted and reflected–reflected outputs, leading to a 50% loss [37] and a key rate of . By contrast, our scheme avoids beam splitters and instead employs two sources, EPS 1 and EPS 2, yielding a combined rate of . Theoretically, this corresponds to a fourfold improvement, [M = (2)/(/2) = 4], in the sifted key rate compared with the conventional BBM92 protocol. Experimentally, we observe more than sixfold enhancement (Fig. 3(a)), which we attribute to non-ideal splitting ratios, inherent bias, and losses in practical beam splitters.
@@
where and denote coincidence counts in the correct and incorrect bases, respectively. As shown in Fig. 3(b), the QBER in our scheme (black dots) increases from 1.03% to 6.44% at a slope of 0.45%/mW, whereas the conventional approach (black squares) rises from 1.8% to 8.04% at 0.55%/mW. These results confirm that the present BBM92 protocol, which exploits the intrinsic spatial randomness of SPDC photons, outperforms the conventional beam-splitter–based scheme by simultaneously achieving higher key rates and lower QBER.
We further verified the relation between entanglement visibility () and QBER () in our scheme, as this serves as a sensitive indicator of eavesdropping in the network. Previous works [39, 40] established that the Bell parameter and QBER are linked through , yielding the critical threshold of for secure key distribution [38]. Since also relates to the interference visibility as , the QBER and visibility are connected by . Experimentally, we varied the quantum interference visibility by tuning the pump power and measured the corresponding QBER using Eq. 1 and Eq. 2. The results are shown in Fig. 4. As expected, the QBER increases with decreasing interference visibility (averaged over four bases), with a slope of 1.78, close to the theoretical value of 2. This agreement confirms that visibility, and thus the derived Bell parameter, provides a reliable diagnostic for security in our current BBM92 protocol. Any tampering by an eavesdropper would reduce entanglement correlations, lowering below 2 and driving the QBER above the critical 11% threshold. Such conditions can be used to terminate the protocol before key generation, ensuring robustness of the system.
@@
We compared the coincidence counts of Alice and Bob in the correct bases (, , , and ) for both our scheme and the conventional BBM92 protocol at a fixed pump power of 3 mW. The results are shown in Fig. 5. As evident from Fig. 5, in the conventional scheme, the coincidence rates are measured to be 3.2, 4.1, 3.1, and 2.7 kHz, respectively, while in our scheme the corresponding values increased to 20.3, 21.6, 21.6, and 21.8 kHz, respectively, showing an order-of-magnitude enhancement in the available bit rate. In addition to the higher count/bit rates, the distribution of coincidence events across the and bases provides a measure of the intrinsic randomness of the encoding. Ideally, the ratio should approach 1:1. For the conventional BBM92 protocol, we find a deviation to 1.3:1, reflecting imperfections of the beam splitter. In contrast, our scheme yields a nearly ideal ratio of 1:1, confirming that the spatially random distribution of SPDC photons ensures unbiased, quantum-mechanical encoding, as also shown in our previous report [34].
@@
To provide a clear comparison between the present BBM92 scheme and the conventional beam-splitter–based BBM92 protocol, we summarize the key performance parameters—including QBER, sifted key rate, error-corrected key rate, privacy-amplified key rate, and the final secure key rate—in Table 1. The calculations employ standard post-processing procedures, with error correction implemented via low-density parity check (LDPC) codes and privacy amplification performed using established algorithms. As shown in Table 1, our spatial-randomness–based scheme consistently outperforms the conventional protocol across all stages of key generation, demonstrating its advantage for high-rate and secure quantum communication.
Parameters | Conventional | Present |
scheme | scheme | |
QBER (%) | 8.04 | 6.44 |
Sifted key rate (Mbps) | 0.47 | 2.90 |
Post EC key rate (Mbps) | 0.45 | 2.76 |
Key rate after PA (Mbps) | 0.15 | 1.02 |
Secure key rate (Mbps) | 0.15 | 1.02 |
4 Conclusions
We demonstrate a new experimental implementation of the efficient and scalable BBM92 QKD protocol by exploiting the quantum mechanical randomness of the spatial distribution of pair photons generated through momentum conservation in the down-conversion process. Using a polarization-entangled photon source with an annular-ring spatial distribution, we derived two entangled photon sources producing pair photons randomly and used for random basis selection, thereby replacing the beam splitter employed in conventional BBM92 schemes. Unlike the conventional approach, this method avoids bit loss due to the imperfections of the physical beam splitters. In a proof-of-concept demonstration over 1 m free space and under identical experimental conditions, the scheme achieved sifted key rate in megabits per second, a 6.4-fold enhancement, consistently lower QBER (7%) across varying pump powers, and a near-ideal encoding balance between logical bits ‘0’s and ‘1’s, thus eliminating encoding bias.
@@ 4 Conclusions
Although the current spatial demultiplexing scheme requires four quantum channels, it can be integrated into fiber-based QKD systems by exploiting energy conservation to generate two wavelength-correlated entangled sources from a single SPDC process and using wavelength demultiplexers in place of beam splitters. Thus, the spatial or wavelength demultiplexing-based randomness establishes a practical route toward high-rate, low-error, and bias-free entanglement-based BBM92 QKD protocol, paving the way for robust and scalable quantum communication networks.
ACKNOWLEDGMENTS
A. K. N., G. P. S, R.G., and G. K. S. acknowledge the support of the Department of Space, Govt. of India. A. K. N. acknowledges funding support for Chanakya - PhD fellowship from the National Mission on Interdisciplinary Cyber-Physical Systems of the Department of Science and Technology, Govt. of India, through the I-HUB Quantum Technology Foundation. G. K. S. acknowledges the support of the Department of Science and Technology, Govt of India, through the Technology Development Program (Project DST/TDT/TDP-03/2022).
@@ ACKNOWLEDGMENTS
A. K. N. developed the experimental setup and performed measurements. G.P.S. participated in experiments, A. K. N., R.G., and C. M. C. participated in data analysis, numerical simulation, and data interpretation. A. K. N. and G. K. S. developed the ideas, and G. K. S led the project. All authors participated in the discussion and contributed to the manuscript writing.
DATA AVAILABILITY

### v2

@@ III Results and discussions
We first characterized EPS 1 and EPS 2 in terms of their quantum performance parameters, including entanglement visibility, Bell parameter, and state fidelity. As both sources are derived from a single entangled photon source, they exhibit nearly identical characteristics. For clarity, we present the representative results of EPS 2 (see green paths in Fig. 1(b)). The measurements, summarized in Fig. 2, were performed at a fixed pump power of 3 mW and crystal temperature of T = 41.5∘C.
@@
We recorded the coincidence counts between photons transmitted through the PBS of both Alice and Bob, with Alice having fast axis of the HWP fixed at , , , and , corresponding to the (black), (blue), (red), and (green) bases, respectively, while Bob continuously rotated the HWP from to in steps. The resulting interference fringes are shown in Fig. 2(a). As shown in Fig. 2(a), the coincidence counts exhibit the expected sinusoidal modulation, with measured fringe visibilities of 97.8, 97.8, 97.7, and 97.7 in the , , , and bases, respectively—well above the quantum threshold of 71 Jabir and Samanta (2017), thereby confirming high-quality entanglement. Furthermore, by analyzing coincidence counts across 16 different combinations of HWP settings of Alice and Bob and applying the standard CHSH inequality formalism Clauser et al. (1969), we obtained a Bell parameter of , clearly violating the classical bound by nearly 109 standard deviations. Further, we carried out quantum state tomography based on coincidence measurements across different polarization projections. The reconstructed real and imaginary parts of the two-qubit density matrix, shown in Fig. 2(b) and 2(c), respectively, confirm the generation of the maximally entangled Bell state with a fidelity of 97. The source brightness of EPS 2 is estimated to be 0.04 MHz/mW/nm. We observed comparable performance parameters for EPS 1. These results establish the suitability of both EPS 1 and EPS 2 for the implementation of the QKD protocol.
@@
With the quantum parameters of the entangled photon sources established, we implemented the BBM92 protocol and benchmarked its performance against the conventional beam-splitter–based approach. Keeping the crystal temperature fixed at C, we measured the sifted key rate and QBER as functions of pump power. The results are shown in Fig. 3. As expected, the sifted key rate (see Fig. 3(a)) calculated by summing all the coincidence counts of same basis detection of two users (here, Alice and Bob), increases with pump power in both cases owing to the higher pair-generation rate. In basis-randomization Protocol #1 (red dots), the sifted key rate increases from 0.29 Mbps to 2.9 Mbps as the pump power is raised from 1 to 13 mW, corresponding to a slope of 0.22 Mbps/mW for a one-meter free-space link. In basis-randomization Protocol #2 (blue dots), the sifted key rate increases from 0.148 Mbps to 1.52 Mbps over the same pump-power range, yielding a slope of 0.11 Mbps/mW. By contrast, the conventional scheme (black dots) exhibits only a modest increase from 0.045 Mbps to 0.47 Mbps, with a slope of 0.035 Mbps/mW. Such a superior performance of our current BBM92 with post detected basis randomization protocols ( 1 and 2) in terms of higher sifted key rate with respect to the conventional BBM92 scheme can be understood as follows. Let us consider one of the derived sources, EPS1. In the absence of a beam splitter, the coincidence rate between Alice and Bob in the same basis (e.g., Z) is . Introducing an ideal 50:50 beam splitter at each photon path of Alice and Bob reduces the useful coincidences to the transmitted–transmitted and reflected–reflected outputs, leading to a 50% loss Neumann et al. (2021) and a key rate of . In Protocol #1, the measurement basis for coincident events is assigned in a correlated manner, thereby eliminating the conventional loss arising from basis mismatch. In contrast, Protocol #2 employs independent basis assignment at Alice and Bob, leading to a 50% loss of coincident events due to basis mismatch, similar to the standard BBM92 protocol. Moreover, our scheme avoids beam splitters and utilizes two entangled photon sources, EPS 1 and EPS 2. Consequently, Protocol #1 and Protocol #2 yield sifted key rates of and , corresponding to fourfold and twofold enhancements, respectively, compared to the conventional BBM92 scheme.
Experimentally, we observe more than sixfold and three fold enhancement (Fig. 3(a)) in the sifted key rate for Protocol 1 and Protocol 2 with respect to conventional BBM92 protocol under similar experimental conditions, which we attribute to non-ideal splitting ratios, inherent bias, and losses in practical beam splitters.
@@
where and denote coincidence counts in the correct and incorrect bases, respectively. As shown in Fig. 3(b), the QBER in our scheme for Protocol #1 (red dots) increases from 1.03% to 6.44% at a slope of 0.45%/mW, and for Protocol #2 (blue dots) increases from 1.43% to 9.28% at a slope of 0.65%/mW, whereas the conventional approach (black dots) rises from 1.8% to 8.04% at 0.55%/mW. These results confirm that the present BBM92 scheme with new basis randomization protocols, which exploits the intrinsic spatial randomness of SPDC photons, outperforms the conventional beam-splitter–based scheme by simultaneously achieving higher key rates and lower QBER.
@@
We computed the coincidence counts between Alice and Bob for both the present scheme and the conventional BBM92 scheme at a fixed pump power of 3 mW. The distribution of coincidence events across the and bases provides a measure of the intrinsic randomness in linear and rectilinear basis encoding. Ideally, this distribution should approach a 1:1 ratio. For the conventional BBM92 scheme, the numbers of encoded bits in the H/V and D/A bases are 74.9 kHz and 58.42 kHz, respectively, yielding a ratio of 1.3:1 and indicating deviations from the ideal case due to beam-splitter imperfections. In contrast, for Protocol #1 the numbers of encoded bits in the H/V and D/A bases are 436.5 kHz and 434 kHz, while for Protocol #2 they are 220 kHz and 220 kHz, respectively, resulting in a nearly ideal 1:1 ratio. These results confirm that the spatially random distribution of SPDC photons enables unbiased, quantum-mechanical encoding, consistent with our previous report Nai et al. (2024).
@@
To provide a clear comparison between the present BBM92 scheme and the conventional beam-splitter–based BBM92 scheme, we summarize the key performance parameters including QBER, sifted key rate, error-corrected key rate, privacy-amplified key rate, and the final secure key rate in Table 1. The calculations employ standard post-processing procedures, with error correction implemented via low-density parity check (LDPC) codes and privacy amplification performed using established algorithms. As shown in Table 1, our spatial-randomness–based scheme with both basis randomization protocols consistently outperforms the conventional protocol across all stages of key generation, demonstrating its advantage for high-rate and secure quantum communication.
| Parameters | Conventional | Present Scheme | |
| Scheme | Protocol #1 | Protocol #2 | |
| Encoding ratio (H/V:D/A) | 1.3:1 | 1:1 | 1:1 |
| QBER (%) | 8.04 | 6.44 | 9.27 |
| Sifted key rate (Mbps) | 0.47 | 2.90 | 1.52 |
| Post EC key rate (Mbps) | 0.45 | 2.76 | 1.44 |
| Key rate after PA (Mbps) | 0.15 | 1.02 | 0.45 |
| Secure key rate (Mbps) | 0.15 | 1.02 | 0.45 |
IV Conclusions
We demonstrate a new experimental implementation of an efficient and scalable BBM92 QKD protocol by exploiting the intrinsic quantum-mechanical randomness of the spatial distribution of photon pairs generated through momentum conservation in the SPDC process. Using a polarization-entangled photon source with an annular spatial-ring distribution, we derive two entangled photon sources that generate photon pairs randomly and used for measurement-basis selection, thereby replacing the beam splitters employed in conventional BBM92 schemes. Unlike the conventional approach, this method avoids bit loss arising from imperfections of physical beam splitters. However, the spatial distinguishability of the photon pairs could, in principle, introduce security vulnerabilities. We mitigate this by performing basis randomization after photon detection using two distinct protocols that exploit the intrinsic timing jitter of the detectors, thereby concealing the basis information from a potential eavesdropper. This post-detection basis assignment effectively emulates stochastic basis selection while avoiding beam-splitter induced losses and bias. The security of the scheme nevertheless relies on the assumptions of trusted detectors, intrinsic and uncontrollable timing jitter, and the operational indistinguishability of the photon sources.
@@ IV Conclusions
In a proof-of-concept demonstration over a 1 m free-space link and under identical experimental conditions, the scheme achieves sifted key rates in the megabits-per-second regime, corresponding to a 6.4-fold enhancement over the conventional BBM92 protocol, consistently lower QBER () across varying pump powers, and a near-ideal encoding balance between linear and rectilinear basis thereby eliminating encoding bias. Although the present spatial demultiplexing scheme requires four quantum channels, it can be adapted to fiber-based QKD systems by exploiting energy conservation to generate two wavelength-correlated entangled sources from a single SPDC process and by using wavelength demultiplexers instead of beam splitters. Spatial- or wavelength-demultiplexing-based randomness thus provides a practical pathway toward high-rate, low-error, and bias-free entanglement-based BBM92 QKD, enabling robust and scalable quantum communication networks.
@@
A. K. N. developed the experimental setup and performed measurements. G. P. S. participated in experiments, A. K. N., R. G., and C. M. C. participated in data analysis, numerical simulation, and data interpretation. A. K. N. and G. K. S. developed the ideas, and G. K. S. led the project. All authors participated in the discussion and contributed to the manuscript writing.
Acknowledgements.
A. K. N., G. P. S., R. G., and G. K. S. acknowledge the support of the Department of Space, Govt. of India. A. K. N. acknowledges funding support for Chanakya - PhD fellowship from the National Mission on Interdisciplinary Cyber-Physical Systems of the Department of Science and Technology, Govt. of India, through the I-HUB Quantum Technology Foundation. G. K. S. acknowledges the support of the Department of Science and Technology, Govt of India, through the Technology Development Program (Project DST/TDT/TDP-03/2022).DATA AVAILABILITY